OpenStudy (anonymous):

PLEASE HELP!!! Create your own equation written in standard form. Determine any point that is a solution and will be on this line. Justify your answer algebraically.

terenzreignz (terenzreignz):

Have you made up an equation?

OpenStudy (anonymous):

yes 6x + 2y = 125

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Okay. Now all we have to do is identify a point on your line. Any point will do. First step here is pick a value for either x or y. Any value will do, although, do have mercy on yourself, and pick a rather simple value.

OpenStudy (anonymous):

how about 2 for x?

OpenStudy (anonymous):

What do I do now?

terenzreignz (terenzreignz):

Replace x with 2. What becomes of the equation.

terenzreignz (terenzreignz):

?

OpenStudy (anonymous):

6(2) +2y = 125?

terenzreignz (terenzreignz):

Yup. Now just solve for y.

OpenStudy (anonymous):

So would the final answer be 56.5?

terenzreignz (terenzreignz):

Peronally, I prefer \(\Large \frac{113}{2}\) But yeah, basically, yes :D So that means the point \[\Large \left(2,\frac{113}2\right)\] is a point on your line.

terenzreignz (terenzreignz):

*Personally

OpenStudy (anonymous):

Thanks :)

terenzreignz (terenzreignz):

By the way, to justify algebraically, just replace x with 2 and y with 56.5 in your made-up equation and show that it does indeed turn into 125. kthxbye XD